Question 11 Rexfleld Company recorded the following transactions for the month of February: sales were $421,400 for the month. Sales prices are determined by a markup on manufacturing cost of 40%. The costs of new inputs to the manufacturing process during the month were $290,000. (a) Calculate the missing values in the schedule below, Materials Work in Process Finished Goods Purchases Beginning Inventory Ending Inventory Direct Materlals Used Direct Labour Manufacturing Overhead (including $105,000 16,000 $8,500 18,800 90,300 18,400 indirect materials of $8,500) 115,600 Transferred to Flnished Goods Cost of Goods Sold
Answer #1

a) Missing amount

Materials Work in process Finished goods
Purchases 105000
Beginning inventory 16000 8500 301000+18400-279700 = 39700
Ending inventory 16000+105000-90300-8500 = 22200 18800 18400
Direct material used 90300
Direct labor 290000-90300-115600) = 84100
Manufacturing overhead (including indirect materials of 8500) 115600
Transferred to finished goods 279700
Cost of goods sold 421400*100/140 = 301000